# Urban Traffic Forecasting with Integrated Travel Time and Data Availability in a Conformal Graph Neural Network Framework

### v1

@@ IV-A Dataset
We used the Ohio ODOT-TCDS dataset with traffic data from 51,777 N-CCS and 396 CCS nodes, recorded every 15 minutes in 2019. It can be observed that the adjacency matrix in Figure 2 is asymmetrical which is due to directional trip times. For example, node 1 connects to nodes 2-9 (1-8 minutes), while the reverse direction takes longer.
@@ IV-B Evaluation Metric and Baselines
To measure and evaluate the performance of different methods, Mean Absolute Errors (MAE), and Root Mean Squared Errors (RMSE) are adopted.
@@
We compare our method with the following baselines:
  • Spatio-Temporal Graph Convolutional Networks (STGCN): Combines graph and temporal convolutions for traffic prediction.
  • Auto-Regressive Integrated Moving Average (ARIMA): Uses past values in a time series for prediction.
  • Historical Average (HA): Predicts based on the average of past traffic data.
  • Feed-Forward Neural Network (FNN): Basic neural network without temporal dependencies.
  • Long Short-Term Memory (LSTM): Captures long-term dependencies in time-series data.
@@ IV-C Results
Table I demonstrates the results for our model and comparison with the baselines on the ODOT-TCDS dataset. Our proposed model shows a reduction of 24% in MAE and 8% in RMSE compared with STGCN.
@@
The results for uncertainty quantification for a selected CCS node are shown in Figure 3. We used a forecast horizon of 15 min with a 24 hours of look-back period. It also shows the associated prediction intervals on actual flow data.
@@ IV-D Traffic Model Calibration and Validation
We used the routesampler tool in SUMO to create a traffic scenario following the methodology in [29] for validation. We took the average values from the upper bound predicted by ACP to mimic a peak-hour scenario for morning hours of weekdays from 7:30 am to 8:30 am. Then, we conducted a Monte-Carlo simulation by varying the departure time of the generated vehicles for 200 runs and measured the VUT travel time parameter to evaluate the model’s performance which was compared against the INRIX probe data [31]. The results are shown in Figure 4.
@@ V CONCLUSIONS
In this study, we propose an approach that uses travel times to define the weights in a GNN’s adjacency matrix, simplifying it while maintaining a physics-like perspective. We also include data from all types of traffic monitoring stations for comprehensive coverage. To address uncertainty, we apply ACP method for better reliability. Finally, we validate the predictions through Monte-Carlo simulation in SUMO, assessing VUT travel time distribution against INRIX data, which showed that the predicted travel times closely matched the actual data.
In order to improve the model’s forecasting ability, future directions for this work will include a variety of urban contexts and the addition of outside variables like traffic accidents and weather fluctuations.

### v2

@@ IV-A Dataset
We used the Ohio ODOT-TCDS data subset with traffic information from 51,777 N-CCS and 396 CCS nodes, recorded every 15 minutes in 2019. Because the trip times between stations are directional, the adjacency matrix in Figure 2 is asymmetrical. For example, node 1 connects to nodes 2-9 (1-8 minutes), while the reverse direction takes much longer.
@@ IV-B Evaluation Metric and Baselines
Mean Absolute Errors (MAE) and Root Mean Squared Errors (RMSE) are adopted to measure and evaluate the performance of different methods.
@@
We compare our method with the following baselines: 1) Spatio-Temporal Graph Convolutional Networks (STGCN) [2]; 2) Auto-Regressive Integrated Moving Average (ARIMA) [28]; 3) Historical Average (HA); 4) Feed-Forward Neural Network (FNN); and 5) LSTM.
@@ IV-C Results
Table I demonstrates the results for our model and comparison with the baselines on the ODOT-TCDS dataset. Our proposed model shows a reduction of 24% in MAE and 8% in RMSE compared with STGCN. We used a forecast horizon of 15 min with a 24 hours of look-back period.
@@
It can be observed that ARIMA and HA perform poorly due to their inability to capture non-linear and dynamic traffic patterns. In contrast, FNN and LSTM improve accuracy by modeling non-linear relationships and temporal dependencies. A much better model, STGCN, further enhances predictions by incorporating spatial features. However, our proposed model surpasses STGCN by integrating travel times and data availability, more effectively capturing complex urban traffic dynamics.
@@
The results for uncertainty quantification for a selected CCS node are shown in Figure 4.
@@ IV-D Traffic Model Calibration and Validation
We utilized the routesampler tool in SUMO following methodology in [29] to generate traffic elements and origin-destination pairs based on vehicle counts. This tool uses linear programming to match real-world traffic counts, adjusting parameters like vehicle departure intervals, traffic volume, driver behavior, and vehicle dynamics for realistic simulations. We extracted the ACP’s upper bound predictions to mimic peak-hour traffic from 7:30 to 8:30 am on weekdays. Then, a Monte-Carlo simulation was conducted by varying the departure time of the generated vehicles for 200 runs and measured the VUT travel time parameter to evaluate the model’s performance which was compared against the INRIX probe data [31], as shown in Figure 5.
@@ V CONCLUSIONS
In this study, we propose an approach that uses travel times to define the weights in a GNN’s adjacency matrix, simplifying it while maintaining a physics-like perspective. We also include data from all types of traffic monitoring stations for comprehensive coverage. To address uncertainty, we apply the ACP method for better reliability. It outperforms existing methods by effectively integrating spatial and temporal data, capturing long-term dependencies, and providing dynamic uncertainty bounds. However, its complexity and data requirements may limit use in areas with sparse data. Finally, we validate the predictions through Monte-Carlo simulation in SUMO, assessing VUT travel time distribution against INRIX probe data, which showed that the predicted travel times closely matched the real-world data.
To improve the model’s forecasting ability, future directions for this work will include a variety of urban contexts and the addition of outside variables like traffic accidents and weather fluctuations.